# Few Batches or Little Memory, But Not Both: Simultaneous Space and Adaptivity Constraints in Stochastic Bandits

### 14 Mar 2026 v1

@@ 5.2 Algorithm Analysis
Each empirical mean is an average of independent $[0,1]$-valued rewards. By Hoeffding’s inequality, for any fixed mean based on $s$ samples we have $\Pr\big(|\widehat{\mu}-\mu|>c(s)\big)\leq\delta$. The total number of empirical means computed is at most the number of batches, which is $O(K\log^{2}\log T)$ by Lemma 18, and hence at most $T^{2}$ for all $T$ large enough. Applying a union bound over all computed means yields $\Pr(E)\geq 1-T^{2}\cdot\delta\geq 1-\frac{1}{T^{2}}$. ∎
@@
During outer iteration $i$, the quantity $\widehat{\mu}_{i}^{\star}$ is initialized to $\widehat{\mu}^{\mathrm{init}}_{i}$ and can only be updated when some challenger $j$ survives through level $i$ and replaces the incumbent, in which case it is set to $\widehat{\mu}_{i,i}(j)$. Therefore $\widehat{\mu}_{i}^{\star}$ is always either the initial incumbent estimate $\widehat{\mu}^{\mathrm{init}}_{i}$ or one of the challenger summaries $\widehat{\mu}_{i,i}(j)$. Both are covered by the event $E$, and both are based on $t_{i}$ samples. Hence $\widehat{\mu}_{i}^{\star}$ is always within $c(t_{i})$ of the true mean of the current incumbent. ∎
@@
The space bound follows from Lemma 16, the schedule and round-budget bounds from Lemma 17, and the batch bound from Lemma 18. It remains to prove the regret bound. We condition on the event $E$ from Lemma 19. Within a fixed outer iteration, $a_{i}^{\star}$ denotes the current incumbent at the moment under discussion; at the end of the iteration it is the final incumbent of that iteration.
The best arm is never eliminated.
Let $a^{\star}\in\arg\max_{j\in[K]}\mu_{j}$ denote a true best arm. Fix an outer iteration $i$, and suppose the algorithm tests $a^{\star}$ at some level $k\leq i$. The elimination condition would be
$\widehat{\mu}_{i,k}(a^{\star})<\widehat{\mu}_{i}^{\star}-2c(t_{k}).$
@@
$\widehat{\mu}_{i,k}(a^{\star})\geq\mu^{\star}-c(t_{k}),\qquad\widehat{\mu}_{i}^{\star}\leq\mu_{a_{i}^{\star}}+c(t_{i})\leq\mu^{\star}+c(t_{i})\leq\mu^{\star}+c(t_{k}),$
@@
$\widehat{\mu}_{i,k}(a^{\star})\geq\widehat{\mu}_{i}^{\star}-2c(t_{k}),$
so the elimination condition can never hold for $a^{\star}$. Hence $a^{\star}$ always survives to level $k=i$ whenever it is tested in outer iteration $i$.
Incumbent true means do not decrease within an outer iteration.
Fix an outer iteration $i$, and suppose a challenger arm $j$ replaces the current incumbent. Immediately before the update, by the update rule,
$\widehat{\mu}_{i,i}(j)\geq\widehat{\mu}_{i}^{\star}+2c(t_{i}).$
@@
$\mu_{j}\geq\widehat{\mu}_{i,i}(j)-c(t_{i})\geq\widehat{\mu}_{i}^{\star}+c(t_{i})\geq\mu_{a_{i}^{\star}},$
where the last inequality uses Lemma 20. Thus every update can only increase the true mean of the incumbent. Consequently, every incumbent that appears during outer iteration $i$ has true mean at least that of the initial incumbent, which is $a^{*}_{i-1}$.
@@
Fix an outer iteration $i$. If $a^{\star}=a_{i,\mathrm{init}}$, then the incumbent is already optimal at the start of the iteration, and the previous paragraph implies that every later incumbent in this iteration is no worse. Thus the final incumbent has gap 0. Now consider the case $a^{\star}\neq a_{i,\mathrm{init}}$. When the algorithm tests $a^{\star}$, we showed that the best arm is never eliminated, so it reaches level $k=i$. At that level, either it updates to $a^{\star}$, in which case the current incumbent becomes $a^{\star}$ and has gap 0, or the update condition fails, i.e.,
$\widehat{\mu}_{i,i}(a^{\star})<\widehat{\mu}_{i}^{\star}+2c(t_{i}).$
In the latter case, under $E$,
$\mu^{\star}-\mu_{a_{i}^{\star}}\leq\bigl(\mu^{\star}-\widehat{\mu}_{i,i}(a^{\star})\bigr)+\bigl(\widehat{\mu}_{i,i}(a^{\star})-\widehat{\mu}_{i}^{\star}\bigr)+\bigl(\widehat{\mu}_{i}^{\star}-\mu_{a_{i}^{\star}}\bigr)<c(t_{i})+2c(t_{i})+c(t_{i})=4c(t_{i}),$
and hence the current incumbent gap is at most $4c(t_{i})$. Moreover, incumbent means do not decrease within an outer iteration, so the final incumbent at the end of outer iteration $i$ also satisfies
@@
Fix an outer iteration $i$, an arm $j$, and a level $k\geq 2$. If arm $j$ reaches level $k$ (i.e., it is not eliminated at level k1k-1), then
$\widehat{\mu}_{i,k-1}(j)\geq\widehat{\mu}_{i}^{\star}-2c(t_{k-1}).$
@@
$\mu_{a_{i}^{\star}}-\mu_{j}\leq\bigl(\mu_{a_{i}^{\star}}-\widehat{\mu}_{i}^{\star}\bigr)+\bigl(\widehat{\mu}_{i}^{\star}-\widehat{\mu}_{i,k-1}(j)\bigr)+\bigl(\widehat{\mu}_{i,k-1}(j)-\mu_{j}\bigr)\leq c(t_{i})+2c(t_{k-1})+c(t_{k-1})\leq 4c(t_{k-1}),$
@@
Since incumbent means do not decrease within an outer iteration, the incumbent present at level k1k-1 cannot be worse than the initial incumbent $a_{i-1}^{\star}$, hence
$\Delta_{a_{i}^{\star}}\leq 4c(t_{k-1}).$
@@
Let $m_{i}(j)$ denote the largest level at which challenger $j$ itself is sampled in outer iteration $i$. If $m_{i}(j)<i$, then the remaining slots $k=m_{i}(j)+1,\dots,i$ are filled by the current incumbent. This decomposition mirrors the predetermined batch layout. The regret therefore has four contributions: (i) incumbent-refresh blocks at the beginning of the outer iterations, (ii) challenger pulls before elimination or replacement, (iii) incumbent-filled tail slots after a challenger is eliminated, and (iv) the final exploitation batch.
@@
(ii) Challenger slots before elimination. The challenger pulls of arm $j$ in outer iteration $i$ occur exactly at levels $k=1,\dots,m_{i}(j)$. The $k=1$ terms contribute at most $O(KLt_{1})$ and are dominated by the final bound. Whenever arm $j$ is pulled at some level $k\geq 2$, it must have survived level k1k-1, and thus (9) applies. Therefore the regret incurred by all challenger pulls at levels $k\geq 2$ is at most
$\sum_{i=1}^{L}\sum_{j\in[K]}\sum_{k=2}^{m_{i}(j)}t_{k}\,\Delta_{j}$ $\leq\sum_{i=1}^{L}\sum_{j\in[K]}\sum_{k=2}^{m_{i}(j)}t_{k}\cdot 8c(t_{k-1})$
$\leq\sum_{i=1}^{L}\sum_{j\in[K]}\sum_{k=2}^{i}t_{k}\cdot 8c(t_{k-1})$
$=O\!\left(\sum_{i=1}^{L}\sum_{j\in[K]}\sum_{k=2}^{i}\frac{t_{k}}{\sqrt{t_{k-1}}}\cdot\sqrt{\log T}\right)$
@@
(iii) Incumbent-filled tail slots. Suppose challenger $j$ is eliminated during outer iteration $i$. Then the remaining slots of this $(i,j)$ comparison are filled by the current incumbent. By the monotonicity established above, every incumbent appearing during outer iteration $i$ is at least as good as the initial incumbent $a_{i-1}^{\star}$. Hence, for $i\geq 2$, each such filler slot incurs regret at most $4c(t_{i-1})$ by (8). Therefore the total regret of these incumbent-filled tail slots is at most
$\sum_{i=2}^{L}\sum_{j\in[K]}\sum_{k=m_{i}(j)+1}^{i}t_{k}\cdot\Delta_{\mathrm{inc},i,j,k}$ $\leq\sum_{i=2}^{L}(K-1)\Bigl(\sum_{k=1}^{i}t_{k}\Bigr)\cdot 4c(t_{i-1})$
$\leq 8(K-1)\sum_{i=2}^{L}t_{i}c(t_{i-1})$
@@
where $\Delta_{\mathrm{inc},i,j,k}$ denotes the gap of the incumbent used to fill that slot. Thus replacing an eliminated challenger by the incumbent does not change the order of the regret bound.
@@ 6 Conclusion
even under adaptive grids. Thus, in the natural word-RAM regime $W=\Theta(\log T)$, the batch complexity of near-minimax learning is essentially linear in $K$ up to polylogarithmic factors.
@@
Several directions remain open. The most immediate algorithmic question is to characterize the optimal batch complexity beyond the logarithmic-memory regime. It would also be interesting to understand whether the same memory–adaptivity bottleneck persists in richer models such as linear, generalized linear, and contextual bandits, where the latent instance information is structured rather than arm-wise independent.

### 30 Mar 2026 v2

@@ 5.2 Algorithm Analysis
Each empirical mean is an average of independent $[0,1]$-valued rewards. By Hoeffding’s inequality, for any fixed mean based on $s$ samples we have $\Pr\big(|\widehat{\mu}-\mu|>c(s)\big)\leq\delta$. The total number of empirical means computed is at most
$L+\sum_{i=1}^{L}(K-1)i=O(KL^{2})=O\!\big(K\log^{2}\log T\big),$
since in outer iteration $i$, each of the K1K-1 challenger positions can produce at most one empirical mean at each of the $i$ levels. Hence this total is at most $T^{2}$ whenever $T\geq cK$ for a sufficiently large universal constant $c$. Applying a union bound over all computed means yields $\Pr(E)\geq 1-T^{2}\cdot\delta\geq 1-\frac{1}{T^{2}}$. ∎
@@
In particular, for every challenger block $C_{i,j}$, its frozen benchmark summary satisfies
$\bigl|\bar{\mu}_{i,j}-\mu_{\bar{a}_{i,j}}\bigr|\leq c(t_{i}).$
@@
During outer iteration $i$, the quantity $\widehat{\mu}_{i}^{\star}$ is initialized to $\widehat{\mu}^{\mathrm{init}}_{i}$ and can only be updated at the end of a challenger block, when some active arm $\tilde{a}_{i,j}$ replaces the incumbent, in which case it is set to $\widehat{\mu}_{i,j,i}(\tilde{a}_{i,j})$. Therefore $\widehat{\mu}_{i}^{\star}$ is always either the initial incumbent estimate $\widehat{\mu}^{\mathrm{init}}_{i}$ or one of the block-level challenger summaries $\widehat{\mu}_{i,j,i}(a)$. Both are covered by the event $E$, and both are based on $t_{i}$ samples. Hence $\widehat{\mu}_{i}^{\star}$ is always within $c(t_{i})$ of the true mean of the current incumbent. Since each frozen benchmark summary $\bar{\mu}_{i,j}$ is just a copy of $\widehat{\mu}_{i}^{\star}$ taken at the start of block $j$, the same bound holds for $(\bar{a}_{i,j},\bar{\mu}_{i,j})$. ∎
@@
The space bound follows from Lemma 15, the schedule and round-budget bounds from Lemma 16, and the batch bound from Lemma 17. It remains to prove the regret bound. We condition on the event $E$ from Lemma 18. Within a fixed outer iteration, $a_{i}^{\star}$ denotes the current incumbent at the moment under discussion; at the end of the iteration it is the final incumbent of that iteration. For each block $C_{i,j}$, let $(\bar{a}_{i,j},\bar{\mu}_{i,j})$ denote the benchmark frozen at the start of that block.
The best arm is never deactivated.
Let $a^{\star}\in\arg\max_{a\in[K]}\mu_{a}$ denote a true best arm. Fix an outer iteration $i$, let $j^{\star}$ be the block containing $a^{\star}$ if $a^{\star}\neq a_{i,\mathrm{init}}$, and suppose the algorithm tests $a^{\star}$ at some level $k\leq i$ within that block. The deactivation condition would be
$\widehat{\mu}_{i,j^{\star},k}(a^{\star})<\bar{\mu}_{i,j^{\star}}-2c(t_{k}).$
@@
$\widehat{\mu}_{i,j^{\star},k}(a^{\star})\geq\mu^{\star}-c(t_{k}),\qquad\bar{\mu}_{i,j^{\star}}\leq\mu_{\bar{a}_{i,j^{\star}}}+c(t_{i})\leq\mu^{\star}+c(t_{i})\leq\mu^{\star}+c(t_{k}),$
@@
$\widehat{\mu}_{i,j^{\star},k}(a^{\star})\geq\bar{\mu}_{i,j^{\star}}-2c(t_{k}),$
so the deactivation condition can never hold for $a^{\star}$. Hence $a^{\star}$ always survives to level $k=i$ whenever it is tested in outer iteration $i$.
Incumbent true means do not decrease across block updates within an outer iteration.
Fix an outer iteration $i$, a block $C_{i,j}$, and suppose its selected arm $\tilde{a}_{i,j}$ replaces the current incumbent. Immediately before the update, by the update rule,
$\widehat{\mu}_{i,j,i}(\tilde{a}_{i,j})\geq\bar{\mu}_{i,j}+2c(t_{i}).$
@@
$\mu_{\tilde{a}_{i,j}}\geq\widehat{\mu}_{i,j,i}(\tilde{a}_{i,j})-c(t_{i})\geq\bar{\mu}_{i,j}+c(t_{i})\geq\mu_{\bar{a}_{i,j}},$
where the last inequality uses Lemma 19. Thus every block update can only increase the true mean of the incumbent. Consequently, every incumbent that appears during outer iteration $i$ has true mean at least that of the initial incumbent, which is $a_{i-1}^{\star}$.
@@
Fix an outer iteration $i$. If $a^{\star}=a_{i,\mathrm{init}}$, then the incumbent is already optimal at the start of the iteration, and the previous paragraph implies that every later incumbent in this iteration is no worse. Thus the final incumbent has gap 0. Now consider the case $a^{\star}\neq a_{i,\mathrm{init}}$. Let $j^{\star}$ be the block containing $a^{\star}$. When the algorithm tests $a^{\star}$ in block $j^{\star}$, we showed that the best arm is never deactivated, so it reaches level $k=i$. At the end of that block, there are two cases.
If the block update succeeds, then the algorithm installs some arm $\tilde{a}_{i,j^{\star}}$ satisfying
$\widehat{\mu}_{i,j^{\star},i}(\tilde{a}_{i,j^{\star}})\geq\widehat{\mu}_{i,j^{\star},i}(a^{\star}),$
because $\tilde{a}_{i,j^{\star}}$ is chosen as the active arm with the largest final empirical mean. Under $E$,
$\mu_{\tilde{a}_{i,j^{\star}}}\geq\widehat{\mu}_{i,j^{\star},i}(\tilde{a}_{i,j^{\star}})-c(t_{i})\geq\widehat{\mu}_{i,j^{\star},i}(a^{\star})-c(t_{i})\geq\mu^{\star}-2c(t_{i}),$
so the updated incumbent has gap at most $2c(t_{i})$.
If the block update fails, then
$\widehat{\mu}_{i,j^{\star},i}(\tilde{a}_{i,j^{\star}})<\bar{\mu}_{i,j^{\star}}+2c(t_{i}).$
Since $a^{\star}$ is still active at level $i$, maximality of $\tilde{a}_{i,j^{\star}}$ implies
$\widehat{\mu}_{i,j^{\star},i}(a^{\star})\leq\widehat{\mu}_{i,j^{\star},i}(\tilde{a}_{i,j^{\star}})<\bar{\mu}_{i,j^{\star}}+2c(t_{i}).$
Therefore, under $E$,
$\mu^{\star}-\mu_{\bar{a}_{i,j^{\star}}}\leq\bigl(\mu^{\star}-\widehat{\mu}_{i,j^{\star},i}(a^{\star})\bigr)+\bigl(\widehat{\mu}_{i,j^{\star},i}(a^{\star})-\bar{\mu}_{i,j^{\star}}\bigr)+\bigl(\bar{\mu}_{i,j^{\star}}-\mu_{\bar{a}_{i,j^{\star}}}\bigr)<c(t_{i})+2c(t_{i})+c(t_{i})=4c(t_{i}),$
and hence the incumbent present after block $j^{\star}$ has gap at most $4c(t_{i})$. Moreover, incumbent means do not decrease within the remaining blocks of outer iteration $i$, so the final incumbent at the end of outer iteration $i$ also satisfies
@@
Fix an outer iteration $i$, a block $C_{i,j}$, an arm $a\in C_{i,j}$, and a level $k\geq 2$. If arm $a$ reaches level $k$ (i.e., it is not deactivated at level k1k-1), then
$\widehat{\mu}_{i,j,k-1}(a)\geq\bar{\mu}_{i,j}-2c(t_{k-1}).$
@@
$\mu_{\bar{a}_{i,j}}-\mu_{a}\leq\bigl(\mu_{\bar{a}_{i,j}}-\bar{\mu}_{i,j}\bigr)+\bigl(\bar{\mu}_{i,j}-\widehat{\mu}_{i,j,k-1}(a)\bigr)+\bigl(\widehat{\mu}_{i,j,k-1}(a)-\mu_{a}\bigr)\leq c(t_{i})+2c(t_{k-1})+c(t_{k-1})\leq 4c(t_{k-1}),$
@@
Since incumbent means do not decrease across block updates within an outer iteration, the block benchmark arm $\bar{a}_{i,j}$ cannot be worse than the initial incumbent $a_{i-1}^{\star}$, hence
$\Delta_{\bar{a}_{i,j}}\leq 4c(t_{k-1}).$
@@
Let $k_{i}(a)$ denote the largest level at which arm $a$ itself is sampled in outer iteration $i$. If $k_{i}(a)<i$, then the remaining slots $k=k_{i}(a)+1,\dots,i$ reserved for arm $a$ are filled by the frozen benchmark of its block. This decomposition mirrors the predetermined batch layout. The regret therefore has four contributions: (i) incumbent-refresh blocks at the beginning of the outer iterations, (ii) challenger pulls before deactivation or replacement, (iii) benchmark-filled tail slots after a challenger is deactivated, and (iv) the final exploitation batch.
@@
(ii) Challenger slots before deactivation. The challenger pulls of arm $a$ in outer iteration $i$ occur exactly at levels $k=1,\dots,k_{i}(a)$ inside the unique block that contains $a$. The $k=1$ terms contribute at most $O(KLt_{1})$ and are dominated by the final bound. Whenever arm $a$ is pulled at some level $k\geq 2$, it must have survived level k1k-1, and thus (11) applies. Therefore the regret incurred by all challenger pulls at levels $k\geq 2$ is at most
$\sum_{i=1}^{L}\sum_{a\in[K]}\sum_{k=2}^{k_{i}(a)}t_{k}\,\Delta_{a}$ $\leq\sum_{i=1}^{L}\sum_{a\in[K]}\sum_{k=2}^{k_{i}(a)}t_{k}\cdot 8c(t_{k-1})$
$\leq\sum_{i=1}^{L}\sum_{a\in[K]}\sum_{k=2}^{i}t_{k}\cdot 8c(t_{k-1})$
$=O\!\left(\sum_{i=1}^{L}\sum_{a\in[K]}\sum_{k=2}^{i}\frac{t_{k}}{\sqrt{t_{k-1}}}\cdot\sqrt{\log T}\right)$
@@
(iii) Benchmark-filled tail slots. Suppose challenger $a$ is deactivated during outer iteration $i$ inside block $C_{i,j}$. Then the remaining slots reserved for this arm are filled by the frozen benchmark arm $\bar{a}_{i,j}$ of that block. By the monotonicity established above, every block benchmark appearing during outer iteration $i$ is at least as good as the initial incumbent $a_{i-1}^{\star}$. Hence, for $i\geq 2$, each such filler slot incurs regret at most $4c(t_{i-1})$ by (10). Therefore the total regret of these benchmark-filled tail slots is at most
$\sum_{i=2}^{L}\sum_{a\in[K]}\sum_{k=k_{i}(a)+1}^{i}t_{k}\cdot\Delta_{\mathrm{inc},i,a,k}$ $\leq\sum_{i=2}^{L}(K-1)\Bigl(\sum_{k=1}^{i}t_{k}\Bigr)\cdot 4c(t_{i-1})$
$\leq 4(1+\sqrt{2})(K-1)\sum_{i=2}^{L}t_{i}c(t_{i-1})$
@@
where $\Delta_{\mathrm{inc},i,a,k}$ denotes the gap of the benchmark arm used to fill that slot. Thus replacing a deactivated challenger by the block benchmark does not change the order of the regret bound.
@@ 6 Conclusion
even under adaptive grids, and our static-grid upper bound shows that this dependence on $K/W$ is achievable up to polylogarithmic factors. Thus the batch complexity of near-minimax learning is governed by the memory–adaptivity trade-off captured by $K/W$.
@@
Several directions remain open. The most immediate algorithmic question is to sharpen the remaining polylogarithmic gap in the $K/W$ trade-off. It would also be interesting to understand whether the same memory–adaptivity bottleneck persists in richer models such as linear, generalized linear, and contextual bandits, where the latent instance information is structured rather than arm-wise independent.
Appendix A Deferred Measurability Proofs
A.1 Localized change of measure: formal setup and proofs
This subsection collects the measurable-space formulation and the truncation-based proofs used in Section 4.1.
Formal setup.
Write $\mathbb{E}_{0},\mathbb{E}_{1}$ for the corresponding expectations.
The algorithm is a non-anticipating policy: for each $t\geq 1$,
$A_{t}=\rho_{t}\bigl(U,A_{1},R_{1},\dots,A_{t-1},R_{t-1}\bigr)\in[K],$
for some measurable map $\rho_{t}$. The observed reward is
$\mathcal{F}_{t}=\sigma(U,A_{1},R_{1},\dots,A_{t},R_{t}).$
For each integer $n\geq 0$, define the prefix sigma-field
Lemma 20 (Budget-event measurability).
For every event $E\in\mathcal{F}_{T}$ and every integer $n\geq 0$,
$E\cap\{N_{j}(T)\leq n\}\in\mathcal{G}_{n}.$
Proof.
and
By construction, the truncated transcript
and the count $\widetilde{N}_{j}^{(n)}(T)$ are $\mathcal{G}_{n}$-measurable.
$\{N_{j}(T)\leq n\}=\{\widetilde{N}_{j}^{(n)}(T)\leq n\}\in\mathcal{G}_{n}.$
Finally, let
$H_{T}:=(U,A_{1},R_{1},\dots,A_{T},R_{T}).$
$E=\{H_{T}\in\Gamma_{E}\}.$
and the right-hand side is $\mathcal{G}_{n}$-measurable. ∎
A.2 Formal proof that the boundary transcript determines the sampling profile
Proof of Lemma 13.
$N_{i}^{(b)}:=\sum_{t=t_{b-1}+1}^{t_{b}}\mathbb{I}\{A_{t}=i\},$
and let
$N^{(b)}:=(N_{i}^{(b)})_{i\in[K]}.$
$N^{(b)}=c_{b}(M_{t_{b-1}},U,\tau)\qquad\text{a.s.}$
$Y_{i}:=\mathbb{I}\Bigl\{\sum_{b=1}^{B}c_{b,i}(M_{t_{b-1}},U,\tau)\geq n\Bigr\},\qquad i\in[K].$
Since $M_{0}$ is fixed, this proves that $Y$ is a deterministic function of
$(U,\tau,M_{t_{1}},\dots,M_{t_{B-1}}).$
$(U,M_{t_{1}},\dots,M_{t_{B-1}}),$